# A Microscopic Traffic Flow Model on Network with Destination-Aware V2V Communications and Rational Decision-Making

### v1

@@ 3 The models with V2V communications
Communication range $\mathcal{R}$ 150 $[0,+\infty)$ m
@@
Memory $\delta m$ +$\infty$+ ∞ $[0,+\infty)$ min
@@ 3.1 Test 2: spreading of information
[…]
@@
Fig. 5a shows the impact of the cascade effect. As expected, removing the cascade effect slows down the spreading of the information. Also, disabling the cascade effect, not only $K_{\mathcal{N}}$ is lower, but it never saturates. Fig. 5b shows the sensitivity with respect to the parameter $\delta c$. Peaks in correspondence of communications are clearly visible. In between, $K_{\mathcal{N}}$ remains constant or decreases as some cars become inactive. Fig. 5c shows the sensitivity with respect to the parameter $\mathcal{R}$. One can see that if $\mathcal{R}$ is low, $K_{\mathcal{N}}$ does not saturate even if the cascade effect is enabled. Fig. 5d shows the sensitivity with respect to the parameter $N_{\textsc{c}}$. As expected, $K_{\mathcal{N}}$ is proportional to $N_{\textsc{c}}$, and the first time $K_{\mathcal{N}}$ saturates is inversely proportional to it. This parameter seems to be the one with the most impact on the information spreading. Fig. 5e shows the sensitivity with respect to the parameter $N_{\textsc{j}}$. In this case we observe the opposite behavior with respect to the previous case, i.e. $K_{\mathcal{N}}$ is inversely proportional to $N_{\textsc{j}}$, while first time of saturation is proportional to it. Finally, Fig. 5f shows the sensitivity with respect to the parameter $\delta m$. If cars have no memory the number of known cars equals the cars within the distance $\mathcal{R}$ at any given time. For low or mid-size values of $\delta m$, the function $K_{\mathcal{N}}$ does not saturate.
@@ 3.2.1 The V2V-RUE model
The model we propose is based on two main assumptions: 1) at each rendez-vous, cars share their current position and destination only; 2) at each fixed time $\bar{t}$, each car is able to nowcast111 nowcast is a term often used in the engineering literature which merges the words ‘now’ and ‘forecast’, meaning a prediction from the past to the current time (now). the positions of the known cars by means of a parallel fictitious simulation. In the parallel simulation done by car c, we assume that the network is populated only by cars in $\mathcal{K}_{\textsc{c}}$, and that they move along a minimum-time path, computed, in turn, employing the road weights $w_{\textsc{c}}(\bar{t},\cdot)$, i.e. the weights based on the c’s knowledge only.
@@
    positions of all cars $\textsc{c}^{\prime}\in\mathcal{K}_{\textsc{c}}(\bar{t})$ are nowcast, in a imaginary fictitious simulation, from the last rendez-vous time $\hat{t}$ to time $\bar{t}$ using (1)-(2), assuming that only cars in $\mathcal{K}_{\textsc{c}}(\bar{t})$ are present on the network. For doing this, a temporary $\textsc{nextroad}^{\sharp}_{\textsc{c}^{\prime}}(s,\cdot)$, $s\in[\hat{t},\bar{t})$ is computed by (4) using $w_{\textsc{c}}(s,\cdot)$, $s\in[\hat{t},\bar{t})$;
@@ 3.2.3 Test 4: a new traffic equilibrium
The interesting point is that neither path can actually be chosen by car c. If we force the car to choose the shortest &\&& congested path $P_{s}$, the car will encounter a large number of other cars (since the path is congested), then the information collected will allow the car to understand that a longer but less congested path is preferable. Conversely, if we force the car to choose the longer &\&& uncongested path $P_{f}$, the limited amount of traffic-related information will not allow the car to recognize that path as optimal, instead the shortest path will be preferred (if car c is not aware of any congestion, it will naturally tend to prefer the shortest path).
@@
Fig. 6 shows a specific system configuration on the $5\times 5$ Manhattan-like network with $N_{\textsc{c}}=50$, for which these three paths (shortest, longer optimal, equilibrium) can be explicitly computed. The simulation is realized by manually assigning origin, destination and whole path to each car on the network. In addition, only for the car of interest, we depict at each passed junction a red arrow indicating the choice that the driver would have made if he had been free to decide his path with V2V-RUE dynamics. If all arrows point toward the road actually chosen, the path is of equilibrium.
@@ 3.3.1 The V2V-DUE model
    positions of all cars $\textsc{c}^{\prime}\in\mathcal{K}_{\textsc{c}}(\bar{t})$ are nowcast, in a imaginary fictitious simulation, from the last rendez-vous time $\hat{t}$ to time $\bar{t}$ using (1)-(2) and $\textsc{nextroad}_{\textsc{c}^{\prime}}(s,\cdot)$, $s\in[\hat{t},\bar{t})$, assuming that only cars in $\mathcal{K}_{\textsc{c}}(\bar{t})$ are present on the network;
@@
    $\textsc{nextroad}_{\textsc{c}}(t,\cdot)$, $\forall t\geqslant\bar{t}$, is recomputed using (6) (in the time interval $[\bar{t},T_{fin}]$);
@@ 3.3.2 Test 5
We have also checked the correctness of the limit behaviors $\mathcal{R}=0$ and $\mathcal{R}=+\infty$. In the first case, the ttt of both V2V-RUE and V2V-DUE always coincides with that of BB. In the second case, the ttt of V2V-RUE always coincides with that of RUE, while for V2V-DUE the situation is more delicate. Whenever DUE exists and forward-backward iterations converge, the ttt of V2V-DUE coincides with that of DUE. If instead DUE is instable and forward-backward iterations do not converge, we are not able to say anything about the convergence of V2V-DUE to DUE.
Finally, let us note that, in some cases, the ttt of V2V-DUE is equal or larger than the ttt of V2V-DUE, see Fig. 7a-7c. Actually, we have also observed that in the setting of Fig. 4b, for some particular initial conditions, also DUE’s ttt is equal or larger than RUE’s ttt. This is not in principle impossible to happen, and we think that the main reason is the great number of equivalently optimal paths in the Manhattan-like network.
@@ 4 Conclusions and future work
Acknowledgements The authors would like to warmly thank Simone Cacace, Davide Vergni, and Alessio Olivierio for their useful comments and suggestions. We also want to send a very special thank to Maurizio Falcone, who introduced us, directly or indirectly, to the beauty of the DPP.
@@ Appendix A Implementation details
  • when method M4 (Remark 3.1) is considered for DUE, the DPP must be slightly modified with respect to the standard version (6). In fact, when one simulates the dynamics of the probe vehicle from the beginning to the end of any road r, one must consider all the possible nextroad’s it can have once it reaches the end of the road (because different nextroad means different nextcar, then different dynamics, as pointed out in the previous comment). As a consequence, the function $w$ depends not only on $t$ and r, but also on any road outgoing from the end of road r. The minimum in (6) must be then extended by covering these additional cases;

### v2

@@ 3 The models with V2V communications
Communication range $\mathcal{R}$ 150 $0,+\infty)$ m
@@
Memory $\delta m$ $\infty$ $0,+\infty)$ min
@@ 3.1 Test 2: spreading of information
$K_{\mathcal{N}}(t):=\frac{1}{N_{a}(t)}\sum_{\text{$\textsc{c}$ is active}}K_{\textsc{c}}(t),$ (9)
@@
Fig. 5a shows the impact of the cascade effect. As expected, removing the cascade effect slows down the spreading of the information. Also, disabling the cascade effect, not only $K_{\mathcal{N}}$ is lower, but it never saturates. Fig. 5b shows the sensitivity with respect to the parameter $\delta c$. Peaks in correspondence of communications are clearly visible. In between, $K_{\mathcal{N}}$ remains constant or decreases as some cars become inactive. Fig. 5c shows the sensitivity with respect to the parameter $\mathcal{R}$. One can see that if $\mathcal{R}$ is small, $K_{\mathcal{N}}$ does not saturate even if the cascade effect is enabled. Fig. 5d shows the sensitivity with respect to the parameter $N_{\textsc{c}}$. As expected, $K_{\mathcal{N}}$ is proportional to $N_{\textsc{c}}$, and the first time $K_{\mathcal{N}}$ saturates is inversely proportional to it. This parameter seems to be the one with the most impact on the information spreading. Fig. 5e shows the sensitivity with respect to the parameter $N_{\textsc{j}}$. In this case we observe the opposite behavior with respect to the previous case, i.e. $K_{\mathcal{N}}$ is inversely proportional to $N_{\textsc{j}}$, while first time of saturation is proportional to it. Finally, Fig. 5f shows the sensitivity with respect to the parameter $\delta m$. If cars have no memory the number of known cars equals the cars within the distance $\mathcal{R}$ at any given time. For small or mid-size values of $\delta m$, the function $K_{\mathcal{N}}$ does not saturate.
@@ 3.2.1 The V2V-RUE model
The model we propose is based on two main assumptions: 1) at each rendez-vous, cars share their current position and destination only; 2) at each fixed time $\bar{t}$, each car is able to nowcast444nowcast is a term often used in the engineering literature which merges the words ‘now’ and ‘forecast’, meaning a prediction from the past to the current time (now). the positions of the known cars by means of a parallel fictitious simulation. In the parallel simulation done by car c, we assume that the network is populated only by cars in $\mathcal{K}_{\textsc{c}}$, and that they move along a minimum-time path, computed, in turn, employing the road weights $w_{\textsc{c}}(\bar{t},\cdot)$, i.e. the weights based on the c’s knowledge only.
@@
    positions of all cars $\textsc{c}^{\prime}\in\mathcal{K}_{\textsc{c}}(\bar{t})$ are nowcast, in an imaginary fictitious simulation, from the last rendez-vous time $\hat{t}$ to time $\bar{t}$ using (1)-(2), assuming that only cars in $\mathcal{K}_{\textsc{c}}(\bar{t})$ are present on the network. For doing this, a temporary $\textsc{nextroad}^{\sharp}_{\textsc{c}^{\prime}}(s,\cdot)$, $s\in[\hat{t},\bar{t})$ is computed by (4) using $w_{\textsc{c}}(s,\cdot)$, $s\in[\hat{t},\bar{t})$;
@@ 3.2.3 Test 4: a new traffic equilibrium
The interesting point is that neither path can actually be chosen by car c. If we force the car to choose the shortest &\& congested path $P_{s}$, the car will encounter a large number of other cars (since the path is congested), then the information collected will allow the car to understand that a longer but less congested path is preferable. Conversely, if we force the car to choose the longer &\& uncongested path $P_{f}$, the limited amount of traffic-related information will not allow the car to recognize that path as optimal, instead the shortest path will be preferred (if car c is not aware of any congestion, it will naturally tend to prefer the shortest path).
@@
Fig. 6555Fig. 6b: www.emilianocristiani.it/attach/paper˙trafficV2Vmicro˙shortest.mp4,
Fig. 6c: www.emilianocristiani.it/attach/paper˙trafficV2Vmicro˙fastest.mp4,
Fig. 6d: www.emilianocristiani.it/attach/paper˙trafficV2Vmicro˙equilibrium.mp4
shows a specific system configuration on the $5\times 5$ Manhattan-like network with $N_{\textsc{c}}=50$, for which these three paths (shortest, longer optimal, equilibrium) can be explicitly computed. The simulation is realized by manually assigning origin, destination and whole path to each car on the network. In addition, only for the car of interest, we depict at each passed junction a red arrow indicating the choice that the driver would have made if he had been free to decide his path with V2V-RUE dynamics. If all arrows point toward the road actually chosen, the path is of equilibrium.
@@ 3.3.1 The V2V-DUE model
    positions of all cars $\textsc{c}^{\prime}\in\mathcal{K}_{\textsc{c}}(\bar{t})$ are nowcast, in an imaginary fictitious simulation, from the last rendez-vous time $\hat{t}$ to time $\bar{t}$ using (1)-(2) and $\textsc{nextroad}_{\textsc{c}^{\prime}}(s,\cdot)$, $s\in[\hat{t},\bar{t})$, assuming that only cars in $\mathcal{K}_{\textsc{c}}(\bar{t})$ are present on the network;
@@
    $\textsc{nextroad}_{\textsc{c}}(t,\cdot)$, $\forall t\geqslant\bar{t}$, is recomputed using (6) (in the time interval $\bar{t},T_{fin}]$);
@@ 3.3.2 Test 5
We have also checked the correctness of the limit behaviors $\mathcal{R}=0$ and $\mathcal{R}=+\infty$. In the first case, the ttt of both V2V-RUE and V2V-DUE always coincides with that of BB. In the second case, the ttt of V2V-RUE always coincides with that of RUE, while for V2V-DUE the situation is more delicate. Whenever DUE exists and forward-backward iterations converge, the ttt of V2V-DUE coincides with that of DUE. If instead DUE is unstable and forward-backward iterations do not converge, we are not able to say anything about the convergence of V2V-DUE to DUE.
Finally, let us note that, in some cases, the ttt of V2V-DUE is equal or larger than the ttt of V2V-RUE, see Fig. 7a-7c. Actually, we have also observed that in the setting of Fig. 4b, for some particular initial conditions, also DUE’s ttt is equal or larger than RUE’s ttt. This is not in principle impossible to happen, and we think that the main reason is the great number of equivalently optimal paths in the Manhattan-like network.
@@ 4 Conclusions and future work
Acknowledgements  We would like to warmly thank Simone Cacace, Davide Vergni, and Alessio Olivierio for their useful comments and suggestions. We also want to send a very special thank to Maurizio Falcone, who introduced us to the beauty of the DPP.
@@
F.L.I. is funded by Sapienza - University of Rome, project “Partial differential equations towards control theory and climate modeling” (project code: RM124190DEC62D0A).
@@ Appendix A Implementation details
  • when method M4 is considered for DUE, the DPP must be slightly modified with respect to the standard version (6). In fact, when one simulates the dynamics of the probe vehicle from the beginning to the end of any road r, one must consider all the possible nextroad’s it can have once it reaches the end of the road (because different nextroad means different nextcar, then different dynamics, as pointed out in the previous comment). As a consequence, the function $w$ depends not only on $t$ and r, but also on any road outgoing from the end of road r. The minimum in (6) must be then extended by covering these additional cases;